3. If P(A) = 0.6, P(B) = 0.55 and P(A \cap B) = 0.2, find P ( (A U B) \ (A \cap B) ).

Please show using diagrams, and what does the symbol " \ " mean?
Answer #1

here the symbol \ means without . P ( (A U B) \ (A \cap B) ) mean that all the events which are in A or B( A u B) but not in A and B (A U B))

hence P ( (A U B) \ (A \cap B) ) =P(A u B)-P(A n B)=P(A)+P(B)-P(A n B)-P(A n B)

=0.6+0.55-0.2-0.2=0.75